OpenStudy (anonymous):

Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant.where (-2,2),(2,-2)

OpenStudy (kc_kennylau):

These questions may help you: (a) What is the slope of the equation? (b) Can you express the equation in point-slope form? (c) Can you convert that to slope-intercept form? Wish you all the best :D

OpenStudy (anonymous):

I am having the problem solving this question

OpenStudy (kc_kennylau):

Where are you stuck, a b or c?

OpenStudy (anonymous):

Kind of all of them.i have trouble this kind of thing

OpenStudy (anonymous):

well let me just say c i guess.

OpenStudy (kc_kennylau):

If you are given two points, you can calculate the slope with the following formula: \[\Large m=\frac{\color{green}{y_2}-\color{blue}{y_1}}{\color{green}{x_2}-\color{blue}{x_1}}\]

OpenStudy (anonymous):

so would the answer be 0 because all of the coordinates are all 2

OpenStudy (kc_kennylau):

You still have to plug the coordinates in, 2 and -2 are different :)

OpenStudy (anonymous):

oh,duh. now i understand.then i would divide the answer i get from subtracting the coordinates

OpenStudy (kc_kennylau):

What have you got now then? :)

OpenStudy (anonymous):

i have -1?

OpenStudy (kc_kennylau):

yep, the slope is -1. Do you remember the point-slope form?

OpenStudy (anonymous):

isn't -1 the point slope form?

OpenStudy (kc_kennylau):

-1 is the slope, not "the equation of the line in point-slope form".

OpenStudy (anonymous):

so how would i find the point-slope?

OpenStudy (kc_kennylau):

You would plug a point in the line named (x1,y1) and the slope named m in the following equation: \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

so i would put, y-(-2)=1(x-(-2)?

OpenStudy (kc_kennylau):

firstly, (a) which point did you choose? (b) what was the slope again?

OpenStudy (anonymous):

the slope was -1 and i choose point -2

OpenStudy (kc_kennylau):

a point has both an x-coordinate and a y-coordinate, one number doesn't make a point :)

OpenStudy (anonymous):

so i should have put -2 and 2?

OpenStudy (kc_kennylau):

yes, -2 as x1 and 2 as y1

OpenStudy (kc_kennylau):

sorry I got to go now, if you want to check, the answer is y=-x, cya :D

OpenStudy (anonymous):

thanks